1. MBR Manufacturing Corp. is considering marketing their new hearing aid in the city of “Big Smoke”. The primary target market for this device is the hearing impaired over the age of 60 years. As she considers the possibility, the VP/Marketing reviews the following data for FY 2021:

Retail price                                                    : $449                                    

Retail margin                                                : 42.5%

Wholesale margin                                        : 27.5%

R & D on hearing aid, FY’s 2019, 2020   : $119,000

Introductory promotional outlays, FY2021: $159,000

MBR’s fixed manufacturing costs             : $155,000 per year (FY 2021)

Variable manufacturing costs/unit             : $95

Retailer’s salesperson’s commission      : 2% of retailer’s selling price

MBR’s sales commission paid                   : 4% of manufacturer’s selling price

Population of Big Smoke”                           : 2,975,000

Proportion of population over 60 years    : 18%

1. How many units must MBR sell in the first year (2021) to break even? Carefully explain, including any assumptions that you make. (6)

2. If 15% of the “over 60” population is hearing impaired, what is MBR’s break-even market share in 2021? (Identify and explain any assumption(s) that are necessary).     (6)

3. Given MBR’s market share (see (c) above), assess MBR’s prospects of breaking even in 2012. (2)
Answer #1

a) Contribution margin = $449 x 0.425 – 95 = 95.825

Break even =( $159,000 + $ 155,000 ) / 95.825 = 3,277 ( rounded ).

b) Overall Population is 2,975,000 so the over 60 populations 18% of that which is 5,35,500. Now only are only using 15% of the overpopulation which totals 80,325 Break Even Unit Volume/Market Served Size=Break Even Market Share 3,277 / 80,325 = .0041 or 4.1%


c) I do not believe he will break even because his market share is so low